beamline experiments
  Beamline Logbook  Logged in as "Eric Dufresne" ELOG logo
Message ID: 58     Entry time: Sat Jun 12 23:45:11 2004
Author: Eric Dufresne 
Type: Fault 
Category: General 
Subject: L5-20 loosing its calibration i.e. its opening 
Hi all, 

today Linda Young and I solved our missing flux problem. 
When we first lined up, we could get up to 1.4 x 10^13 ph/s in 7ID-B
but within less than 24 hours, we lost our flux and could not get it back by
a tweak of the second crystal. The 7ID-B flux went as low as 7 x 10 ^12 ph/s.

Today we found the problem: The L5-20 aperture moved even though their EPICS 
record did not change at all. It was particularly bad in the horizontal
where if I closed the blade physically so than they just barely overlap, the
EPICS opening was 0.162 mm, so with a .5 mm opening, we only had an opening
of 0.338 mm.  It was also noted though on the vertical opening where we
found that when the blade were physically closed, the EPICS record was -.076
mm, so that at a 0.5 mm EPICS opening, we had a 0.576 mm opening. 

So Linda and I recentered the L5-20 slits, recalibrated their openings (I
had done this on June 9, p59 of the log book. I had done this quite
carefully, and the flux was consistent then with our recovered flux now. So
I think the blade moved on their own. There was a power bump 6/12 at 1 am,
and we had two days of bad weather. Perhaps some noise drove the slit away.
Our flux is good again i.e. 1.47 x 10^13 ph/s in top-up hybrid mode. 
The BPM X = 252 um, Y = -183um. Linda had to retweak tonight at 7h40pm, which I
attribute to the monochromator warming up again following all these slit scans.  

Below are some of the scans we did.

Cheers,

Eric 

Attachment 1: vertical center scan with 0.1 mm opening. Note it was not
quite reproducible, as we compare in the figure scan 33 and 34 of file
LYoung-lineup. The FWHM of the scan was 0.426 mm, which seems close to what is
expected. 

Attachment 2: Two repeat scan of the Horizontal Right blade, with zero around
-0.448 mm. The overlap appears not to be perfect but a simple shift of 0.5
microns overlaps the two scans perfectly (not shown). So the Horizontal Right
blade moves reproducibly still. 

Attachment 3: Two repeat scan of the Horizontal Right blade, with good overlap
so the blade moves well as well.

Attachment 4: Scan of the Horizontal center of the L5-20 with an opening of 0.1
mm. The FWHM of the white beam incident on the monochromator is 0.98 mm, 26.5 m
from the source. 
Attachment 1: scan33-34.png  3 kB
Attachment 2: scan36-37.png  5 kB
Attachment 3: scan38-39.png  3 kB
Attachment 4: scan41.png  2 kB
ELOG V2.5.2